# Two liquid states of distinguishable helium-4: the existence of another non-superfluid frozen by heating

### v2

@@ V.2 On state transition
As shown in Sec. III.2, there is no discontinuity in the $H$-$T$ plot and no maximum in the $C_{P}$-$T$ relation even when the LQD-HQD state transitions occur. This denies that the LQD-HQD transitions are the first or second-order phase transitions and supports that they are continuous crossovers. This is in contrast to the fact that the ideal Bose gas exhibits a cusp at the BEC temperature $T_{\rm{C}}$ in the $C_{V}$-$T$ plot [5] and that even the ideal quantum Boltzmann gas possesses a peak as well [74]. For ideal Bose gas, as the number of atoms decreases from the thermodynamic limit ($N\rightarrow\infty$), the cusp of the $C_{V}$-$T$ plot becomes smeared-out [75]. This indicates a finite-size effect; in the $C_{P}$-$T$ relation of the confined 4He, not the $\lambda$-like peak but the blunted shape emerges [66, 76, 77]. However, the lack of maximum in the $C_{P}$-$T$ plot in this study is not because of the size effect. Although the number of atoms in the present simulations is finite ($N=256$), the periodic boundary condition ensures to reproduce the bulk properties [45, 21, 61]. Therefore, the smooth change of $H$ and $V$ with varying temperature in the LQDL-HQDL transition is considered as being a bulk property. This indicates that this transition is continuous crossover, not the thermodynamic transition.
@@
V.3 State metastability and statistical mechanical effect
@@
In Fig. 5, we can see that the data points of He-I, He-II, and solid phase of real 4He lie on a universal line. This line, in other words, denote the $K$-$V$ relation which corresponds to the global minimum of the free energy, since He-I, He-II, and solids phase are all at thermodynamic equilibrium. However, the points of HQDL are extensively scattered on the plane and are downward deviated from the universal line. Therefore, if we introduced the Bosonic permutation to the present system, these deviated points of HQDL would be metastable points. This means that the HQDL can exist only as a metastable state (if it can exist), not as an equilibrium state in reality (i.e., in the regime of Bose statistics).
@@
However, we would mention that, plausibly, HQDL is metastable even in Boltzmann statistics. One main reason is that the HQDL-LQDL transition is continuous without any indication of thermodynamic transition. The LQDL-to-HQDL transition by isobaric cooling is similar to a glass transition in that a metastable state is formed from an equilibrium state without discontinuous change of enthalpy. Both transitions should depend on the production process and hysteresis. The second reason is, being intuitive, that the wide scatter of $K$-$V$ points of HQDL seemingly reflects the ease of emergence of many configurations of stretched flexible necklaces. This can be plausible because the springs of atomic necklaces are remarkably loosed by low temperatures. There should be many local free-energy minima corresponding to necklace configurations in such stretched flexible polymer system.
@@
In fact, we examined another CMD simulation of isothermal decompression of the crystal of distinguishable 4He for 0.1-0.9 K, starting from 100 bar (the pressure at experimental solid phase region)  [^2] M. Tsujimoto and K. Kinugawa, to be submitted.. Even down to 1 bar, the crystal state persisted at the $P$-$T$ conditions where the LQDL or HQDL were produced in the main simulation of this study. We obtained, for example, the enthalpy at 0.5 K and 1 bar, $H=-0.057\pm 0.004$ kJmol-1 for HQDL and $-0.058\pm 0.004$ kJmol-1 for the crystal. Namely, the enthalpy difference is negligible within the standard deviation. Therefore, we cannot say that the crystal is energetically stabler than HQDL. Rather, this crystal state should be regarded as being as metastable as HQDL. Accordingly, both HQDL and the crystal obtained from the decompression must be metastable states under Boltzmann statistics. Every state produced in the experimental He-II region by the Boltzmann-type simulations should plausibly be metastable and correspond to each local free-energy minimum. As a result, during the simulations each state was allowed to exist quasi-stably as it was. If we introduced Bose statistics to this system, these local minima would be converged into a global free-energy minimum, which corresponds to the true thermodynamic equilibrium state, He-II. This should be akin to the gelation of sol, which is equivalent to the emergence of BEC by polymerizing the necklaces [79]. It seems to be difficult to judge which metastable state is the stablest and at the thermodynamic equilibrium in the Boltzmann regime, because energetic difference is quite negligible and estimation of entropy is burdensome. Further examination of the metastability would be a future subject. Regardless of metastability, in the present investigation, we found that one metastable state called the HQDL below 25 bar succeeded in non-freezing down to 0.08 K; this is the conclusion. Finally, we would comment on the stability of distinguishable 4He described in the enlightening report by Boninsegni et al [6]. They reported that at 0.5 K and 24.3 cm3mol-1 the crystal state persisted in distinguishable 4He, whereas it was melted into superfluid state under Bose statistics [6]. We conceive that this retention of the crystal state in the Boltzmann regime may be the same as we described here, whereas Bose statistics certainly plays an indispensable role in emerging the non-freezing superfluid as the authors claim.
@@ VI CONCLUSIONS
In the present study, the path integral CMD simulation for distinguishable 4He obeying Boltzmann statistics was conducted in the range of $0.08\leq{T}\leq{3.3}$ K and $1\leq{P}\leq{500}$ bar. The $P$-$T$ thermodynamic conditions were strictly controlled by means of the isothermal-isobaric CMD method to provide computational results with high quantitativeness. On the basis of the isothermal compression simulations, we revealed the state diagram of this ideal model system on a $P$-$T$ plane. The conclusions are summarized as follows:
@@
(4) The presumed melting line on the $P$-$T$ plane is downward convex and have a minimum point at $T\simeq{0.4}$ K and $P\simeq{25}$ bar, from which the LQDL-HQDL boundary line emanates, as an indication of liquid polyamorphism. Since there is no $C_{P}$ maximum at this transition, the LQDL-HQDL transition is not a thermodynamic phase transition but a continuous crossover. Observable structural change by the transition is smooth and continuous. Thus, the existence of these two liquids is not conventional classical polyamorphism but the quantum polyamorphism characterized by the change of atomic quantum dispersion. The temperature dependence of quantum wavelength and of its expansion factor remarkably changes on the occasion of the LQDL-HQDL transition.
@@
(6) A part of $P$-$T$ region of HQDL overlaps with that of the real non-superfluid states (the LBEC and the BEC-like low-entropy state) of 4He confined in nanopores in meso-porous materials. Since the finite system size suppresses the Bosonic permutation effect in confined 4He, the confined system should be more Boltzmann-like (or suppressed Bosonic) than the bulk. Accordingly, we speculate that HQDL has relevance to these confined system including suppressed Bosonic correlation. We also anticipate that some HQDL-like property can be observed in experiments targeted on the confined systems.
(7) The $P$-$T$ region of the HQDL causing inverse freezing almost overlaps with the reported supersolid region neighboring to the Kim-Chan state boundary. We therefore anticipate that some HQDL-like properties can be also observed in real 4He solid which is put in this $P$-$T$ range.
@@
This is the first investigation that revealed the state diagram, the existence of liquid polyamorphism, and inverse freezing in distinguishable 4He without inclusion of Bosonic statistical effect. Although the present Boltzman-type 4He model is unreal for lower temperatures, the relevance to the real 4He systems could be speculated. This relevance is expected for not bulk systems but some systems of finite size. In real 4He existing in narrow domains such as confined space, dislocations, and disordered domains in a crystal, the long-range Bosonic permutation effect should partially be suppressed, so that the complete BEC does not emerge. We speculate that such semi-Boltzmann-like system are fairly akin to the distinguishable 4He model investigated in this work, so that some HQDL-like properties would likely be observed there in the range of $P\geq 25$ bar and $T\leq 0.4$ K; this is the $P$-$T$ region where HQDL and real 4He solid phase overlap on the state diagram, and is the reported supersolid region.
Finally, we should comment that there remains an unsolved question, “Does distinguishable 4He transit to another state when cooled below 0.1 K?” It would be intriguing to explore whether another state would emerge before reaching absolute zero temperature solely by NQE without disturbance from Bose statistics. Even though liquid exists at $T=0$, it does not violate the third law of thermodynamics so far as the system is in a single state without degeneration [5, 82]. It is also unrevealed whether even more “classical” systems such as Ne and Ar can melt at extremely low temperatures solely by the NQE when the temperature is extremely lowered toward zero temperature. The universality of non-freezing property and state emergence induced by NQE is still an open question.
@@ Appendix A TEST OF FREEZING BY HEATING
In Appendix, we provide the results of another CMD simulation to test the inverse freezing. In Fig. 1, we can see that an isobaric heating of the HQDL at 40 bar, starting from 0.2 K toward 0.3 K, may induce solidification or inverse freezing (freezing by heating). To check this, starting from the end configuration of the production run at 0.2 K and 40 bar, we conducted three isobaric CMD runs. One of them was carried out keeping the temperature at 0.2 K, while the other runs were the process of raising the set temperature to 0.3 or to 0.4 K.
The results of the MD time evolution of $\lambda_{\rm{quantum}}$ are shown in Fig. 14. In this figure, $\lambda_{\rm{quantum}}$ retains at about 3.5 Å when the temperature is kept at 0.2 K. In contrast, it starts to be shortened right after the temperature switching to 0.3 or 0.4 K at $t_{\rm{MD}}=5$ ps. The quantum wavelength finally resulted in $\lambda_{\rm{quantum}}=1.7-1.8$, which is even shorter than 2.2 Å at 0.3 K (LQDA) shown in Fig. 7.
@@
Figure 14: The MD time evolution of quantum wavelength of 4He atoms by isobaric heating test at 40 bar. The graphs show the results of three CMD runs with different control of temperature: (1) keeping the temperature at 0.2 K throughout the run; (2) raising temperature to 0.3 K at $t_{\rm{MD}}=5$ ps; (3) raising temperature to 0.4 K at $t_{\rm{MD}}=5$ ps.
@@
This test simulation showed that the inverse freezing (endothermic freezing) of HQDL was actually caused by isobaric heating from 0.2 to 0.3-0.4 K at 40 bar. This transition $P$-$T$ agrees with the HQDL-LQDA boundary line drawn in the state diagram (Fig. 1). Thus, the present test yielded the same conclusion as we saw in Sec. IV. Although the $P$-$T$ condition of the state boundary in Fig. 1 cannot completely be free from possible hysteresis of production simulation, the present agreement ensures that the HQDL existing in the upward convex region in Fig. 1 certainly freezes on the occasion of crossing over the drawn HQDL-LQDA boundary line. We note again that this line almost overlaps with the Kim-Chan boundary.
@@
Figure SA-3: The $P$-$T$ conditions set for the CMD simulations. The black dots denote 357 conditions (21 temperatures times 17 pressures). The shown phases are experimental ones. Before conducting isothermal compression simulations, the isobaric cooling was conducted at 1 bar to generate the equilibrium state at each temperature at 1 bar.
@@
Figure SA-4: Schematic representation of CMD procedure. (a) Isobaric cooling run at 1 bar conducted prior to starting the isothermal compression; (b) Isothermal compression run started from each temperature point at 1 bar.
@@
Figure SA-5: The $xy$-projected snapshots of the configurations of the 4He atomic necklaces and centroids at 40 bar. Green spheres and white ones denote the centroids and the beads, respectively. The drawn scales of the centroids and the beads are arbitrary.
@@
Figure SA-6: The plot of enthalpy versus temperature over the whole pressure range.
@@
Figure SA-7: The plot of kinetic energy versus temperature.
@@
Figure SA-8: The plot of internal energy versus temperature.
@@
Figure SA-9: The plot of potential energy versus temperature.
@@
Figure SA-10: The isobaric heat capacity obtained from the polynomial curve fitting of enthalpy-temperature relation at 1 bar.
@@
Figure SA-11: The temperature dependence of apparent self-diffusion coefficient above 200 bar. The colors of the symbols are the same as those displayed in Fig. 1 in the main text: LQDL (blue), LQDA (cyan), HQDL (red), and HQDA (magenta). Green horizontal dashed line denotes the threshold between liquids and glasses.
@@
Figure SA-12: The quantumness parameter dependence of apparent self-diffusion coefficient below 100 bar. The colors of the symbols are the same as those displayed in Fig. 1 in the main text: LQDL (blue), LQDA (cyan), HQDL (red), and HQDA (magenta). Green horizontal dashed line denotes the threshold between liquids and glasses.

### v3

@@ V.1 On state transition
As shown in Sec. III.2, there is no discontinuity in the $H$-$T$ plot and no maximum in the $C_{P}$-$T$ relation even when the LQD-HQD state transitions occur. This denies that the LQD-HQD transitions are the first- or second-order phase transitions and supports that they are continuous crossovers. This is in contrast to the fact that the ideal Bose gas exhibits a cusp at the BEC temperature $T_{\rm{C}}$ in the $C_{V}$-$T$ plot [6] and that even the ideal quantum Boltzmann gas possesses a peak as well [65]. For ideal Bose gas, as the number of atoms decreases from the thermodynamic limit ($N\rightarrow\infty$), the cusp of the $C_{V}$-$T$ plot becomes smeared-out [66]. This indicates a finite-size effect; in the $C_{P}$-$T$ relation of the confined 4He, not the $\lambda$-like peak but the blunted shape emerges [67, 68, 69]. However, the lack of maximum in the $C_{P}$-$T$ plot in this study is not because of the size effect. Although the number of atoms in the present simulations is finite ($N=256$), the periodic boundary condition ensures to reproduce the bulk properties [46, 21, 64]. Therefore, the smooth change of $H$ and $V$ with varying temperature in the LQDL-HQDL transition is considered as being a bulk property. This indicates that this transition is continuous crossover, not the thermodynamic transition.
@@
V.2 State metastability and statistical mechanical effect
@@
In Fig. 5, we can see that the data points of He-I, He-II, and solid phase of real 4He lie on a universal line. This line, in other words, denotes the $K$-$V$ relation which corresponds to the global minimum of the free energy, since He-I, He-II, and solids phase are all at thermodynamic equilibrium. However, the points of HQDL are extensively scattered on the plane and are downward deviated from the universal line. Therefore, if we introduced the Bosonic permutation to the present system, these deviated points of HQDL would be metastable points. This means that the HQDL can exist only as a metastable state (if it can exist), not as an equilibrium state in reality (i.e., in the regime of Bose statistics).
@@
However, we would mention that, plausibly, HQDL is metastable even in Boltzmann statistics. One main reason is that the HQDL-LQDL transition by varying temperature is continuous without any indication of thermodynamic transition. The LQDL-to-HQDL transition by isobaric cooling should be similar to a glass transition in that a metastable state is formed from an equilibrium state without discontinuous change of enthalpy. Both transitions should depend on the production process and hysteresis. The second reason is, being intuitive, that the wide scatter of $K$-$V$ points of HQDL seemingly reflects the ease of emergence of many configurations of stretched flexible necklaces. This can be plausible because the springs of atomic necklaces are remarkably loosed by low temperatures. There should be many local free-energy minima corresponding to necklace configurations in such stretched flexible polymer system.
@@
In fact, we examined another CMD simulation of isothermal decompression of the crystal of distinguishable 4He for 0.1-0.9 K, starting from 100 bar (the pressure at experimental solid phase region)  [^2] M. Tsujimoto and K. Kinugawa, to be submitted.. Even down to 1 bar, the crystal state persisted at the $P$-$T$ conditions where the LQDL or HQDL were produced in the main simulation of this study. We obtained, for example, the enthalpy at 0.5 K and 1 bar, $H=-0.057\pm 0.004$ kJmol-1 for HQDL and $-0.058\pm 0.004$ kJmol-1 for the crystal. Namely, the enthalpy difference is negligible within the standard deviation. Therefore, we cannot say that the crystal is energetically stabler than HQDL. Rather, this crystal state should be regarded as being as metastable as HQDL. Accordingly, both HQDL and the crystal obtained from the decompression must be metastable states under Boltzmann statistics. Every state produced in the experimental He-II region by the Boltzmann-type simulations should plausibly be metastable and correspond to each local free-energy minimum. As a result, during the simulations each state was allowed to exist quasi-stably as it was. If we introduced Bose statistics to this system, these local minima would be converged into a global free-energy minimum, which corresponds to the true thermodynamic equilibrium state, He-II. This should be akin to the gelation of sol, which is equivalent to the emergence of BEC by polymerizing the necklaces [71]. It seems to be difficult to judge which metastable state is the stablest and at the thermodynamic equilibrium in the Boltzmann regime, because energetic difference is quite negligible and estimation of entropy is burdensome. Further examination of the metastability would be a future subject. Regardless of the discussion of metastability, it is important that one metastable state called HQDL below 25 bar succeeded in non-freezing down to 0.1 K in the present investigation.
As in our isothermal decompression test, the retention of crystal state at 0.5 K and 24.3 cm3mol-1 was reported for distinguishable 4He [5]. This can also be attributed to the above-mentioned metastability of each state in distinguishable 4He. The retention of the crystalline states, observed in our and their simulations, seems to be nontrivial, considering that distinguishable 4He should be liquid at $T=0$ as described in Sec. I. However, simple extrapolation from the zero-temperature limit does not predict what state or liquid polymorph emerges at a given $P$-$T$ point for $T>0$.
@@
In Appendix B, we compared the present state diagram with the experimental results of confined sub-nano scale systems and supersolid of real 4He.
@@ VI CONCLUSIONS
In the present study, the path integral CMD simulation for distinguishable 4He obeying Boltzmann statistics was conducted in the range of $0.08\leq{T}\leq{3.3}$ K and $1\leq{P}\leq{500}$ bar. The $P$-$T$ thermodynamic conditions were strictly controlled by means of the isothermal-isobaric CMD method to provide computational results with high quantitativeness. On the basis of the isothermal compression simulations, we revealed the state diagram of this ideal model system on a $P$-$T$ plane ($0.1\leq{T}\leq{3.3}$ K and $1\leq{P}\leq{500}$ bar). The conclusions are summarized as follows:
@@
(4) The presumed melting line on the $P$-$T$ plane is downward convex and has a minimum point at $T\simeq{0.4}$ K and $P\simeq{25}$ bar, from which the LQDL-HQDL boundary line emanates, as an indication of liquid polyamorphism. Since there is no $C_{P}$ maximum at this transition, the LQDL-HQDL transition is not a thermodynamic phase transition but a continuous crossover. Observable structural change by the transition is smooth and continuous. Thus, the existence of these two liquids is not conventional classical polyamorphism but the quantum polyamorphism characterized by the change of atomic quantum dispersion. The temperature dependence of quantum wavelength and of its expansion factor remarkably changes on the occasion of the LQDL-HQDL transition.
@@
Finally, in Appendix B, the state diagram of this study was compared with the experimental reports of confined systems and supersolid of real 4He. A part of $P$-$T$ region of HQDL overlaps with that of the real non-superfluid states (the LBEC and the BEC-like low-entropy state) of 4He confined in nanopores in meso-porous materials, which should be more Boltzmann-like (or suppressed Bosonic) than the bulk. Further, the $P$-$T$ region of the HQDL causing inverse freezing almost overlaps with the reported supersolid region neighboring to the Kim-Chan state boundary. These comparisons provide a future subject as to a question whether HQDL has some essential relevance to these real systems.
This is the first investigation that revealed the state diagram, the existence of liquid polyamorphism, and inverse freezing in distinguishable 4He without inclusion of Bosonic statistical effect. As a future subject, it is intriguing to investigate whether the sole NQE would generate another liquid polymorph when the system is cooled down to far lower temperatures such as 1 mK before reaching $T=0$, at which distinguishable 4He should be a liquid. It is also unrevealed whether even more “classical” systems such as Ne and Ar can melt at extremely low temperatures solely by the NQE when the temperature is extremely lowered. The universality of quantum liquid polyamorphism induced by NQE is still an open question.
@@ Appendix A TEST OF FREEZING BY HEATING
In Appendix A, we provide the results of another CMD simulation to test the inverse freezing. In Fig. 1, we can see that an isobaric heating of the HQDL at 40 bar, starting from 0.2 K toward 0.3 K, may induce solidification or inverse freezing (freezing by heating). To check this, starting from the end configuration of the production run at 0.2 K and 40 bar, we conducted three isobaric CMD runs. One of them was carried out keeping the temperature at 0.2 K, while the other runs were the process of raising the set temperature to 0.3 or to 0.4 K.
The results of the MD time evolution of $\lambda_{\rm{quantum}}$ are shown in Fig. 13. In this figure, $\lambda_{\rm{quantum}}$ retains about 3.5 Å when the temperature is kept at 0.2 K. In contrast, it starts to be shortened right after the temperature switching to 0.3 or 0.4 K at $t_{\rm{MD}}=5$ ps. The quantum wavelength finally resulted in $\lambda_{\rm{quantum}}=1.7-1.8$, which is even shorter than 2.2 Å at 0.3 K (LQDA) shown in Fig. 7.
@@
Figure 13: The MD time evolution of quantum wavelength of 4He atoms by isobaric heating test at 40 bar. The graphs show the results of three CMD runs with different control of temperature: (1) keeping the temperature at 0.2 K throughout the run; (2) raising temperature to 0.3 K at $t_{\rm{MD}}=5$ ps (pointed by a magenta arrow); (3) raising temperature to 0.4 K at $t_{\rm{MD}}=5$ ps (pointed by a magenta arrow).
@@
This test simulation showed that the inverse freezing (endothermic freezing) of HQDL was actually caused by isobaric heating from 0.2 to 0.3-0.4 K at 40 bar. This transition $P$-$T$ agrees with the HQDL-LQDA boundary line drawn in the state diagram (Fig. 1). Thus, the present test yielded the same conclusion as we saw in Sec. IV. Although the $P$-$T$ condition of the state boundary in Fig. 1 cannot completely be free from possible hysteresis of production process, the present agreement ensures that the HQDL existing in the upward convex region in Fig. 1 certainly freezes on the occasion of crossing over the drawn HQDL-LQDA boundary line. This line almost overlaps with the Kim-Chan boundary, as described in Appendix B.2.
@@
Figure 15: The comparison of state diagram of distinguishable 4He and confined real 4He systems. LQDL (blue): low quantum dispersion liquid; HQDL (red): high quantum dispersion liquid; LQDA (cyan): low quantum dispersion amorphous solid (metastable); HQDA (magenta): high quantum dispersion amorphous solid (metastable). The Kim-Chan phase boundary is cited from Ref. 32. The phase boundary lines of real 4He doped in Gelsil (25 Å) [76], folded sheets meso-porous materials (FSM) (28 Å) [77], and MCM-41 (core diameter 47 Å) [78], are drawn together. “$T_{\rm{C}}$, FSM”: superfluid onset temperature of the FSM system; “$T_{\rm{B}}$, FSM”: rapid growth temperature of the superfluid fraction of the FSM system [77]; “$T_{\rm{C}}$, Gelsil”: superfluid transition temperature of the Gelsil system; “$T_{\rm{B}}$, Gelsil”: localized BEC (LBEC) temperature of the Gelsil system [76]. Experimentally, the BEC-like low-entropy state exists in gray slant stripe region [77], while the LBEC exists in gray polka-dot region [76].
Appendix B STATE DIAGRAM COMPARISON WITH NON-SUPERFLUID STATES OF REAL SYSTEMS
In Appendix B, withholding speculative interpretation, the present state diagram is compared with that of two types of real systems, (1) several confined 4He systems and (2) reported supersolid.
B.1 Comparison with confined 4He
Recent experiments targeting confined 4He revealed that the non-superfluid states, unlike He-I and He-II, exist at temperatures lower than the ${\rm{\lambda}}$-transition line of the bulk system [79, 67, 80, 76, 81, 82, 83, 77, 84, 78, 85]. Under such confined conditions, the off-diagonal long-range Bosonic correlation is hindered by the wall of glass substrate pores, so that Bosonic permutation effect should partially be suppressed [81, 76]. Figure 15 shows the comparison of our state diagram with that of confined 4He in three kinds of substrates. Shirahama and co-workers revealed that the localized BEC (LBEC) state of 4He confined in Gelsil exists in the region between the boundaries of “$T_{\rm{B}}$, Gelsil” and “$T_{\rm{C}}$, Gelsil” [79, 67, 80, 76, 81]. For 4He confined in the folded sheets meso-porous materials (FSM) (28 Å), Taniguchi et al. reported that the state at temperatures between “$T_{\rm{B}}$, FSM” and “$T_{\rm{C}}$, FSM” was not He-I nor He-II but a BEC-like state with low entropy [77]. In this figure, the experimental state boundaries have the same negative slopes as the HQDL-LQDA boundary line. Some part of the $P$-$T$ regions of the LBEC and the low-entropy BEC-like states overlaps with the upward convex region ($18\lesssim{P}\lesssim{35}$ bar and $T\lesssim{0.35}$ K) of the HQDL causing inverse freezing.
B.2 Comparison with reported supersolid 4He region
In Figs. 1 and 15, the HQDL-LQDA boundary emanating from the point at $T=0.3$ K and $P=26$ bar traces well the Kim-Chan boundary line [32]. This boundary was interpreted as an evidence of supersolid which might exist at lower-temperature side of the line, on the basis of the detection of non-classical moment of inertia by the torsional oscillator experiment [32]. However, nowadays the existence of defect-less crystalline supersolid phase is almost denied, while this anomaly is attributed to some disorder due to dislocation, grain boundaries, glassy domains, and 3He impurities [86]. In such a narrow domain unlike a bulk, the long-range atomic permutation should partially be suppressed so that the BEC should not completely be fulfilled. The HQDL has the upward convex $P$-$T$ region at $T\lesssim 0.3$ K and $P\gtrsim 25$ bar, while it causes such inverse freezing as shown in Appendix A. This $P$-$T$ region of HQDL coincides with the supersolid region.
@@ B.2 Comparison with reported supersolid 4He region
Figure SA-3: The temperature dependence of quantum wavelength and expansion factor with different Trotter numbers at 1 bar. (a) quantum wavelength; (b) expansion factor. The color of the symbols is the same as that displayed in FIG. 7 and FIG. 8 in the main text: LQDL (blue) and HQDL (red). The definition of these properties is given in Sec. III D. The plotted results of $N_{\rm{b}}=500$ and $N=256$ involve the 1 bar data shown in FIG. 7 and FIG. 8(a) in the main text.
@@
Figure SA-4: The bead-bead radial distribution fucntion with different Trotter numbers at 1 bar. (a) comparison between $N_{\rm{b}}=500$ and $N_{\rm{b}}=1500$. (b) comparison between $N_{\rm{b}}=500$ and $N_{\rm{b}}=2000$. The plotted results of $N_{\rm{b}}=500$ and $N=256$ involve the 1 bar data shown in FIG. 6(b) in the main text. We can see that the transition between the LQDL and HQDL occurs between 0.7 and 0.5 K for all $N_{\rm{b}}$ conditions.
@@
Figure SA-5: The $P$-$T$ conditions set for the CMD simulations. The black dots denote 357 conditions (21 temperatures times 17 pressures). The shown phases are experimental ones. Before conducting isothermal compression simulations, the isobaric cooling was conducted at 1 bar to generate the equilibrium state at each temperature at 1 bar.
@@
Figure SA-6: Schematic representation of CMD procedure. (a) Isobaric cooling run at 1 bar conducted prior to starting the isothermal compression; (b) Isothermal compression run started from each temperature point at 1 bar.
@@
Figure SA-7: The $xy$-projected snapshots of the configurations of the 4He atomic necklaces and centroids at 40 bar. Green spheres and white ones denote the centroids and the beads, respectively. The drawn scales of the centroids and the beads are arbitrary.
@@
Figure SA-8: The plot of enthalpy versus temperature over the whole pressure range.
@@
Figure SA-9: The plot of kinetic energy versus temperature.
@@
Figure SA-10: The plot of internal energy versus temperature.
@@
Figure SA-11: The plot of potential energy versus temperature.
@@
Figure SA-12: The isobaric heat capacity obtained from the polynomial curve fitting of enthalpy-temperature relation at 1 bar.
{midpage}
Figure SA-13: The temperature dependence of apparent self-diffusion coefficient above 200 bar. The colors of the symbols are the same as those displayed in Fig. 1 in the main text: LQDL (blue), LQDA (cyan), HQDL (red), and HQDA (magenta). Green horizontal dashed line denotes the threshold between liquids and glasses.
{midpage}
Figure SA-14: The quantumness parameter dependence of apparent self-diffusion coefficient below 100 bar. The colors of the symbols are the same as those displayed in Fig. 1 in the main text: LQDL (blue), LQDA (cyan), HQDL (red), and HQDA (magenta). Green horizontal dashed line denotes the threshold between liquids and glasses.